OpenStudy (anonymous):

Find the number of permutations in the word “geometry”. Find the number of ways to listen to 4 CDs from a selection of 8 CDs. How many different 12-member juries can be chosen from a pool of 25 people? A pizza place offers 5 different cheeses and 10 different toppings. In how many ways can a pizza be made with 2 cheese and 4 toppings? A delicatessen offers 4 different breads, 4 cheeses, and 6 different meats. In how many ways can a sandwich be made with 1 bread,1 cheese and 2 meats?

OpenStudy (anonymous):

one question at a time please

OpenStudy (anonymous):

1.8!/2! 2.8*7*6*5 3.25!/13! 4.5C2*10C4 5.4C1*4C1*6C2

OpenStudy (kc_kennylau):

1. \(\Large\frac{8!}{2!}\) 2. \(\Large\mbox P^8_4=\frac{8!}{4!}\) 3. \(\Large\mbox C^{25}_{12}=\frac{25!}{13!12!}\) 4. \(\Large\mbox C^5_2\times\mbox C^{10}_4=\frac{5!}{3!2!}\cdot\frac{10!}{6!4!}\) 5. \(\Large\mbox C^4_1\times\mbox C^4_1\times\mbox C^6_2=\frac{4!}{3!1!}\cdot\frac{4!}{3!1!}\cdot\frac{6!}{4!2!}\)